### v2

@@ 4 A priori estimates and stability for BSDEs
$\leq\sqrt{2}\left\|\int_{0}^{\cdot}(\tilde{\mathbb{E}}|K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big)|^{2})^{1/2}dW_{s}\right\|_{BMO}\int_{u}^{t_{k+1}}(\mathbb{E}|\Delta Z_{s}|^{2})^{1/2}ds$
$\leq\varepsilon\int_{u}^{t_{k+1}}(\mathbb{E}|\Delta Z_{s}|^{2})^{1/2}ds\qquad u\in I_{k}.$
Taking $\mathbb{E}^{L}[\cdot\,|\,\mathcal{F}_{u}]$ in the local BSDE yields
$|\Delta Y_{u}|$ $\leq\mathbb{E}^{L}[|\Delta Y_{t_{k+1}}|\,|\,\mathcal{F}_{u}]+\varepsilon\int_{u}^{t_{k+1}}\big(\mathbb{E}|\Delta Z_{s}|^{2}\big)^{1/2}ds,\qquad u\in I_{k}.$
Iterating this estimate over the intervals $I_{k+1},I_{k+2},\dots,I_{N-1}$ and using $\Delta Y_{T}=\Delta\xi$, we obtain
$|\Delta Y_{u}|\leq\mathbb{E}^{L}[|\Delta\xi|\,|\,\mathcal{F}_{u}]+\varepsilon\sqrt{T}\,\left(\int_{u}^{T}\mathbb{E}|\Delta Z_{s}|^{2}\,ds\right)^{1/2},\qquad u\in[0,T].$
By BMO energy inequality (kazamaki2006BMO, Sec. 2.1) and Corollary 4.5, the last term is bounded by a constant depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$ since:
$\int_{u}^{T}\Big({\mathbb{E}}|Z_{s}^{1}|^{2}+{\mathbb{E}}|Z_{s}^{2}|^{2}\Big)\,ds\;$ $\leq\;\Big\|\int_{0}^{\cdot}Z_{s}^{1}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}+\Big\|\int_{0}^{\cdot}Z_{s}^{2}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}.$
Therefore, letting $\varepsilon\to 0$ yields
$|\Delta Y_{u}|\leq\mathbb{E}^{L}[|\Delta\xi|\,|\,\mathcal{F}_{u}]=(\Lambda_{u})^{-1}\,\mathbb{E}[\Lambda_{T}|\Delta\xi|\,|\,\mathcal{F}_{u}],\qquad u\in[0,T].$ (4.22)
Now fix $q\in(1,q^{*})$ and let $p$ be its conjugate exponent. By the reverse Hölder inequality and conditional Hölder’s inequality,
$|\Delta Y_{u}|^{p}$ $\leq(\Lambda_{u})^{-p}\,\mathbb{E}[(\Lambda_{T})^{q}\,|\,\mathcal{F}_{u}]^{p/q}\,\mathbb{E}[|\Delta\xi|^{p}\,|\,\mathcal{F}_{u}]$
$\leq(C_{q}^{*})^{p/q}\,\mathbb{E}[|\Delta\xi|^{p}\,|\,\mathcal{F}_{u}],\qquad u\in[0,T].$
By Doob’s maximal inequality,
$\mathbb{E}\Big(\sup_{u\in[0,T]}|\Delta Y_{u}|^{p}\Big)\ \leq\ C_{Y}(p)\,\mathbb{E}|\Delta\xi|^{p},\qquad p\geq p^{*},$ (4.23)
where $p^{*}$ is the conjugate exponent of $q^{*}$ and $C_{Y}(p):=\Big(\frac{p}{p-1}\Big)^{p}(C_{q}^{*})^{p/q}$. They both depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$ as we discussed in the Preliminary step.
$\mathbb{H}^{p}$–control of $\Delta Z$. Fix $p>p^{*}$, and applying Itô’s formula to $|\Delta Y|^{2}$ on $[0,T]$. we obtain
$|\Delta Y_{0}|^{2}+\!\int_{0}^{T}|\Delta Z_{s}|^{2}\,ds=|\Delta\xi|^{2}+2\int_{0}^{T}\Delta Y_{s}\,\widetilde{\mathbb{E}}[\Delta g(s,Z_{s},\tilde{Z}_{s})]\,ds-2\int_{0}^{T}\Delta Y_{s}\,Z_{s}\cdot dW_{s}$
Using the same Fefferman/Cauchy–Schwarz argument as above and $\|\Delta Y\|_{\mathbb{S}^{\infty}}<\infty$, we obtain
$2\int_{0}^{T}\Delta Y_{s}\,\widetilde{{\mathbb{E}}}\!\big[\Delta g(s,Z_{s},\tilde{Z}_{s})\big]\,ds$ $\leq 2K_{z}\int_{0}^{T}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)\,|\Delta Y_{s}|\,|\Delta Z_{s}|\,ds$
$\quad+2K_{\tilde{z}}\int_{0}^{T}|\Delta Y_{s}|\,\widetilde{{\mathbb{E}}}\!\Big[(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|)\,|\Delta\tilde{Z}_{s}|\Big]\,ds$
$\leq\sup_{t\in[0,T]}|\Delta Y_{t}|^{2}\Big(2K_{z}^{2}\int_{0}^{T}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)^{2}\,ds\Big)+\frac{1}{2}\int_{0}^{T}|\Delta Z_{s}|^{2}\,ds$
$\quad+2K_{\tilde{z}}\|\Delta Y\|_{\mathbb{S}^{\infty}}\int_{0}^{T}\widetilde{{\mathbb{E}}}\!\Big[(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|)\,|\Delta\tilde{Z}_{s}|\Big]\,ds.$
The last term vanishes in the limit $\varepsilon\to 0$ by the same slicing argument as in (4.21). Next, let $K:=\int_{0}^{T}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)^{2}\,ds$. Then
${\mathbb{E}}\Big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}\,K^{p/2}\Big]\leq\Big({\mathbb{E}}\big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{2p}\big]\Big)^{1/2}\Big({\mathbb{E}}\big[K^{p}\big]\Big)^{1/2}.$ (4.24)
If $\int_{0}^{\cdot}Z_{s}^{i}\cdot dW_{s}$ is a BMO martingale ($i=1,2$), the BMO energy inequality (see (kazamaki2006BMO, Sec. 2.1)) gives
${\mathbb{E}}\Big[\Big(\int_{0}^{T}|Z_{s}^{i}|^{2}ds\Big)^{p}\Big]\leq p!\,\Big\|\int_{0}^{\cdot}Z_{s}^{i}\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2p}.$
Therefore,
${\mathbb{E}}[K^{p}]\leq 3^{p-1}\Big((3T)^{p}+3^{p}p!\,\Big\|\int_{0}^{t}Z_{s}^{1}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2p}+3^{p}p!\,\Big\|\int_{0}^{t}Z_{s}^{2}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2p}\Big)=:C_{\mathrm{energy}}(p).$ (4.25)
Lastly, with the help of BDG inequality and Cauchy–Schwarz, we have
@@
${\mathbb{E}}\Big|\int_{0}^{T}(\Delta Y_{s})\Delta Z_{s}\cdot dW_{s}\Big|^{p/2}$ $\leq C_{\mathrm{BDG}}\,{\mathbb{E}}\Big[\Big(\int_{0}^{T}|\Delta Y_{s}|^{2}|\Delta Z_{s}|^{2}\,ds\Big)^{p/4}\Big]$
$\leq C_{\mathrm{BDG}}\,\Big({\mathbb{E}}\big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}\big]\Big)^{1/2}\Big({\mathbb{E}}\big[(\int_{0}^{T}|\Delta Z_{s}|^{2}\,ds)^{p/2}\big]\Big)^{1/2}.$ (4.26)
Putting (4.24), (4.25) and (4) together, we obtain
${\mathbb{E}}\Big[\Big(\int_{0}^{T}|\Delta Z_{s}|^{2}ds\Big)^{p/2}\Big]$ $\leq 2^{\frac{3p}{2}-1}{\mathbb{E}}[|\Delta\xi|^{p}]+2^{\frac{3p}{2}-1}(2K_{z}^{2})^{p/2}\sqrt{C_{\mathrm{energy}}(p)}\Big({\mathbb{E}}[\sup_{t\in[0,T]}|\Delta Y_{t}|^{2p}]\Big)^{1/2}$
$\quad+2^{4p-4}(C_{\mathrm{BDG}})^{2}\,{\mathbb{E}}[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}],$
From previous step (4.23), there exist constants $C_{Y}(p),C_{Y}(2p)$ such that
$\mathbb{E}\Big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}\Big]\leq C_{Y}(p)\,\mathbb{E}\big[|\Delta\xi|^{p}\big],\qquad\Big(\mathbb{E}\Big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{2p}\Big]\Big)^{1/2}\leq\sqrt{C_{Y}(2p)}\,\Big(\mathbb{E}\big[|\Delta\xi|^{2p}\big]\Big)^{1/2}.$
Substituting these bounds into the previous inequality, we obtain for all $p\geq p^{*}$,
@@
$\mathbb{E}\Big[\Big(\int_{0}^{T}|\Delta Z_{s}|^{2}ds\Big)^{p/2}\Big]$ $\leq 2^{\frac{3p}{2}-1}\mathbb{E}\big[|\Delta\xi|^{p}\big]+2^{4p-4}(C_{\mathrm{BDG}})^{2}\,C_{Y}(p)\,\mathbb{E}\big[|\Delta\xi|^{p}\big]$
$\quad+2^{\frac{3p}{2}-1}(2K_{z}^{2})^{p/2}\sqrt{C_{\mathrm{energy}}(p)}\,\sqrt{C_{Y}(2p)}\,\Big(\mathbb{E}\big[|\Delta\xi|^{2p}\big]\Big)^{1/2}.$
Finally, since $\mathbb{E}|\Delta\xi|^{p}\leq\big(\mathbb{E}|\Delta\xi|^{2p}\big)^{1/2}$, we may absorb the first two terms into the last one and conclude that
$\mathbb{E}\Big[\Big(\int_{0}^{T}|\Delta Z_{s}|^{2}ds\Big)^{p/2}\Big]$ $\leq C_{Z}(p)\,\Big(\mathbb{E}\big[|\Delta\xi|^{2p}\big]\Big)^{1/2},$
where
$C_{Z}(p):=2^{\frac{3p}{2}-1}+2^{4p-4}(C_{\mathrm{BDG}})^{2}\,C_{Y}(p)+2^{\frac{3p}{2}-1}(2K_{z}^{2})^{p/2}\sqrt{C_{\mathrm{energy}}(p)}\,\sqrt{C_{Y}(2p)}.$
Combing the stability estimate (4.23) completes the proof. ∎
@@ 5 Proof: existence and uniqueness
In this section, we establish the main result for the mean-field BSDE (2.6) under Assumptions $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q})$. The proof proceeds in two main steps. First, we prove the existence of a solution when the terminal condition $\xi$ is Malliavin differentiable with bounded derivative. This is achieved in Lemma 5.1 via a truncation and regularization argument. With the help of the estimates for Lipschitz BSDEs derived in Proposition 3.1, this allows us to apply the classical result from Buckdahn2009MFBSDE.
In the second step, we extend this existence result to the case of general bounded terminal conditions by approximating $\xi$ with a sequence of Malliavin differentiable random variables. For each approximating BSDE, existence is guaranteed by Lemma 5.1, and we invoke the stability result from Theorem 4.6 to show that the sequence of solutions is Cauchy in suitable norms. The limit of this sequence is then shown to solve the original mean-field BSDE (2.6). Uniqueness follows directly from the same stability result. The full argument is completed in the proof of the main theorem below.
@@
Let $(\rm H_{0})$} and $(\rm H_{q})$} holds and suppose $\xi$ is Malliavin differentiable with bounded derivative $\left\|D\xi\right\|_{\mathbb{S}^{\infty}}<\infty$, then the mean-field BSDE admits a solution in $\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}.$
@@
We define a sequence of driver functions $(\bar{g}_{k})_{k\in\mathbb{N}}$ as follows:
$\bar{g}_{k}(t,z,\tilde{z}):=g\left(t,\frac{\left|z\right|\wedge k}{\left|z\right|}z,\frac{\left|\tilde{z}\right|\wedge k}{\left|\tilde{z}\right|}\tilde{z}\right),\quad k\in\mathbb{N}.$
Since $g$ satisfies Assumption $(\rm H_{q})$}, it is straightforward to verify that the new driver $\bar{g}_{k}$ is Lipschitz, with
@@
for all $k\in\mathbb{N}$ and $(t,z,z^{\prime},\tilde{z},\tilde{z}^{\prime})\in[0,T]\times[\mathbb{R}^{d}\times\mathbb{R}^{d}]^{2}$. Thus, the corresponding BSDE with driver $\bar{g}_{k}$ satisfies $(\rm H_{0})$} and $(H_{\mathrm{lip}})$. Then, the mean-field BSDE
@@
admits a unique solution $(Y^{k},Z^{k})\in\mathbb{S}^{2}\times\mathbb{H}^{2}$ (see Theorem 3.1 in Buckdahn2009MFBSDE). Moreover, Proposition 3.1 ensures that for each $k\in\mathbb{N}$, the pair $(Y^{k},Z^{k})$ belongs to $\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$, and the process $Z^{k}$ satisfies a uniform bound independent of the Lipschitz constant of the driver with respect to $z$ and $\tilde{z}$. In particular, there exists a constant $C_{Z}>0$ such that $\left\|Z^{k}\right\|_{\mathbb{S}^{\infty}}\leq C_{Z}$ for all $k\in\mathbb{N}$. Since $(\tilde{Y}^{k},\tilde{Z}^{k})$ are independent copies of $(Y^{k},Z^{k})$, the same uniform bound holds, we can choose $k^{*}\in\mathbb{N}$ large enough so that
$k^{*}>\sup_{k\in\mathbb{N}}\left\|Z^{k}\right\|_{\mathbb{S}^{\infty}}.$
For such $k^{*}$, the truncation is inactive and
$\bar{g}_{k^{*}}(t,Z_{s}^{k^{*}},\tilde{Z}_{s}^{k^{*}})=g(t,Z_{s}^{k^{*}},\tilde{Z}_{s}^{k^{*}}),\quad 0\leq s\leq T.$
Hence, the pair $(Y^{k^{*}},Z^{k^{*}})$ indeed solves the mean-field BSDE (2.6), which completes the proof. ∎
We are now in a position to state the main result of this section. For the reader’s convenience, we restate the main theorem below. To complete the proof of following theorem, it remains to extend the existence result from Malliavin differentiable terminal conditions to the general bounded case. Uniqueness then follows directly from the stability result established in Theorem 4.6.
@@
Let Assumptions $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q})$ hold. Then, there exists a unique solution $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the BSDE (2.6).
@@
We begin by approximating the bounded terminal condition $\xi$ with a sequence of Malliavin differentiable random variables of the form $\xi^{n}=\Phi^{n}(W_{t_{1}},\dots,W_{t_{n}})$, where each $\Phi^{n}$ is a bounded continuous function. Since $\xi$ is bounded, the sequence $(\xi^{n})_{n\in\mathbb{N}}$ can be chosen such that $\xi^{n}$ converges to $\xi$ in in any $L^{p}$ space for $p\geq 1$.
To ensure that the approximating BSDEs satisfy Assumption $(\mathrm{H}_{0})$, we truncate each $\Phi^{n}$ by setting $\Phi^{n}:=\Phi^{n}\wedge\left\|\xi\right\|_{L^{\infty}}$. Hence, each $\xi^{n}$ remains bounded, and $(\xi^{n},g)$ satisfies $(\mathrm{H}_{0})$ for all $n\in\mathbb{N}$. By Lemma 5.1, each approximating BSDE with terminal condition $\xi^{n}$ admits a solution $(Y^{n},Z^{n})\in\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$, satisfying

### v3

@@ 4 A priori estimates and stability for BSDEs
$\left\|\Delta Z\cdot W\right\|_{\mathrm{BMO}_{[u,v]}(\mathbb{P})}\leq\frac{C}{2\sqrt{2}K_{\tilde{z}}}\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{[u,v]}(\tilde{\mathbb{P}})}$.
In addition, since $t\mapsto\left(\mathbb{E}\int_{0}^{t}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)^{2}ds\right)^{1/2}$ is uniformly continuous, for any $\varepsilon>0$, there exists $\delta>0$ such that $v-u<\delta$ implies
$\left(\mathbb{E}\int_{u}^{v}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)^{2}ds\right)^{1/2}<\frac{\varepsilon}{C}.$ (4.25)
As a result, (4) becomes
@@
$|\Delta Y_{t}|^{2}+\mathbb{E}^{L}_{t}\int_{t}^{v}|\Delta Z_{s}|^{2}ds$ $\leq\mathbb{E}^{L}_{t}|\Delta Y_{v}|^{2}+\varepsilon\mathbb{E}^{L}_{t}\sup_{s\in[u,v]}|\Delta Y_{s}|\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{[u,v]}(\tilde{\mathbb{P}})}$
@@
$\leq\left\|\Delta Y_{v}\right\|^{2}_{L^{\infty}}+\frac{\varepsilon}{2}\left\|\sup_{s\in[u,v]}|\Delta Y_{s}|\right\|^{2}_{L^{\infty}}+\frac{\varepsilon}{2}\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{[u,v]}(\tilde{\mathbb{P}})}^{2}.$
By taking $\varepsilon=1/2$, by the similar argument in the proof of previous theorem, we have
$\left\|\sup_{s\in[u,v]}|\Delta Y_{s}|^{2}\right\|_{L^{\infty}}+\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{[u,v]}(\tilde{\mathbb{P}})}^{2}\leq 4\left\|\Delta Y_{v}\right\|^{2}_{L^{\infty}}.$
If we partition $[0,T]$ into intervals of size less than $\delta$ and apply the above estimates iteratively from $T$ to 0, we obtain the estimate
$\left\|\sup_{s\in[0,T]}|\Delta Y_{s}|^{2}\right\|_{L^{\infty}}+\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}(\tilde{\mathbb{P}})}^{2}\leq N4^{N}\left\|\Delta\xi\right\|_{L^{\infty}}^{2},$
where $N$ is the number of subintervals in the partition. Another application of Lemma 2.2 proves the claim. ∎
@@ 5 Proof: existence and uniqueness
In this section, we prove the existence and uniqueness results for the mean-field BSDE (2.6) in the two quadratic regimes introduced earlier. Under Assumptions $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{unif}})$}, we establish Theorem 5.3 for terminal conditions of the form $\xi=\phi(W_{t_{1}},\ldots,W_{t_{n}})$ with $\phi$ bounded and $\alpha$-Hölder continuous. The generalization to Theorem 2.6 can be achieved by measure-change and time-change. Under Assumptions $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{sep}})$}, we establish Theorem 5.4 for terminal conditions of the form $\xi=\phi(W_{[0,T]})$ with $\phi$ bounded and uniformly continuous in the sup norm.
The proofs follow the same general scheme, but with different approximation methods adapted to the corresponding terminal regularity. We first prove in Lemma 5.1 that if the terminal condition $\xi$ is Malliavin differentiable with uniformly bounded derivative, then (2.6) admits a solution in $\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$. The argument is based on truncating the quadratic generator to obtain Lipschitz generator.
Next, we approximate the desired terminal condition by a sequence $(\xi^{n})$ with bounded Malliavin derivatives. In the unified case $(\mathrm{H}_{q}^{\mathrm{unif}})$, we use mollification of the Hölder function $\phi$ on $(\mathbb{R}^{d})^{n}$ and verify, via Lemma 5.2, the uniform quadratic-variation bounds required in Theorem 4.6. In the separable case $(\mathrm{H}_{q}^{\mathrm{sep}})$, we use the Moreau–Yosida approximation on path space, which yields Lipschitz functionals with bounded Malliavin derivatives. For each $n$, Lemma 5.1 provides a solution $(Y^{n},Z^{n})$, and the corresponding stability theorem (Theorem 4.6 under $(\mathrm{H}_{q}^{\mathrm{unif}})$ or Theorem 4.7 under $(\mathrm{H}_{q}^{\mathrm{sep}})$) implies that $(Y^{n},Z^{n})$ is Cauchy in the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$. Passing to the limit yields a solution to (2.6), and uniqueness follows from the same stability result.
@@
Assume $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\mathrm{unif}})$} hold, and suppose $\xi$ is Malliavin differentiable with bounded derivative $\left\|D\xi\right\|_{\mathbb{S}^{\infty}}<\infty$. Then the mean-field BSDE admits a solution in $\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$.
@@
For $k\in\mathbb{N}$, define the truncation map $\pi_{k}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ by
$\pi_{k}(z):=\frac{|z|\wedge k}{|z|}\,z\,\mathbf{1}_{\{z\neq 0\}},\qquad\pi_{k}(0):=0,$
and define a sequence of drivers $(\bar{g}_{k})_{k\in\mathbb{N}}$ by
$\bar{g}_{k}(t,z,\tilde{z}):=g\big(t,\pi_{k}(z),\pi_{k}(\tilde{z})\big),\qquad(t,z,\tilde{z})\in[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{d}.$
Note that $|\pi_{k}(z)|\leq k$ for all $z$, and $\pi_{k}$ is 11–Lipschitz:
$|\pi_{k}(z)-\pi_{k}(z^{\prime})|\leq|z-z^{\prime}|,\qquad\forall z,z^{\prime}\in\mathbb{R}^{d}.$ (5.26)
@@
[…]
Since $|\pi_{k}(\cdot)|\leq k$, the common factor is bounded by $1+4k$, and by (5.26),
[…]
Thus, $\bar{g}_{k}$ satisfies $(\mathrm{H}_{0})$ and $(\mathrm{H}_{\mathrm{lip}})$. Consequently, the mean-field BSDE with driver $\bar{g}_{k}$,
@@
admits a unique solution $(Y^{k},Z^{k})\in\mathbb{S}^{2}\times\mathbb{H}^{2}$ (see Theorem 3.1 in Buckdahn et al. (2009b)). Moreover, Proposition 3.1 ensures that for each $k\in\mathbb{N}$, $(Y^{k},Z^{k})\in\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$. In particular, there exists a constant $C_{Z}>0$ such that
$\|Z^{k}\|_{\mathbb{S}^{\infty}}\leq C_{Z},\qquad\forall k\in\mathbb{N}.$
Since $(\tilde{Y}^{k},\tilde{Z}^{k})$ are independent copies of $(Y^{k},Z^{k})$, the same bound holds for $\tilde{Z}^{k}$. Choose $k^{*}\in\mathbb{N}$ such that $k^{*}>C_{Z}$. Then for this $k^{*}$ the truncation is inactive along the solution and therefore
$\bar{g}_{k^{*}}(s,Z_{s}^{k^{*}},\tilde{Z}_{s}^{k^{*}})=g(s,Z_{s}^{k^{*}},\tilde{Z}_{s}^{k^{*}}),\qquad 0\leq s\leq T.$
Hence, $(Y^{k^{*}},Z^{k^{*}})$ solves the original mean-field BSDE (2.6), which completes the proof. ∎
For the following results, let us define the time partition $0=t_{0}\leq t_{1}\leq t_{2}\leq\cdots\leq t_{n}=T$.
Lemma 5.2.
Consider a bounded function $\phi:(\mathbb{R}^{d})^{n}\to\mathbb{R}$ which is $\alpha$-Hölder continuous, that is, there exists $K>0$ and $\alpha\in(0,1]$ such that
$|\phi(x)-\phi(x^{\prime})|\leq K\max_{i=1,2,\cdots,n}|x_{i}-x^{\prime}_{i}|^{\alpha},$
for any $x,x^{\prime}\in(\mathbb{R}^{d})^{n}$. Then, for $H$ satisfying
$\phi(W_{t_{1}},W_{t_{2}},\cdots,W_{t_{n}})=\mathbb{E}\phi(W_{t_{1}},W_{t_{2}},\cdots,W_{t_{n}})+\int_{0}^{T}H_{s}dW_{s},$
there exists a constant $C$ such that $\int_{0}^{T}|H_{s}|^{2}ds\leq C$ a.s.
Proof.
Let us first prove for the case when $\phi(\omega)=\psi(\omega_{T})$ where $\psi$ is $\alpha$-Hölder continuous function from $\mathbb{R}^{d}$ to $\mathbb{R}$. Then, there exists a function $u(t,x)$ such that $u(T,x)=\psi(x)$ and $\partial_{t}u+\frac{1}{2}\partial^{2}_{xx}u=0$. Then, we have
$\psi(W_{T})=u(0,0)+\int_{0}^{T}\partial_{x}u(t,W_{t})dW_{t}\text{ and }u(t,x)=\mathbb{E}\left[\psi(W_{T})|W_{t}=x\right].$
Since
$u(t,x)=\mathbb{E}\psi(x+W_{T}-W_{t})=\int\psi(x+y)f_{T-t}(y)dy=\int\psi(u)f_{T-t}(u-x)du,$
when $f_{T-t}$ is the PDF of N(0,Tt)N(0,T-t), for $Z\sim N(0,T-t)$, we have
$\partial_{x}u(t,x)$ $=\int\psi(u)\frac{u-x}{T-t}f_{T-t}(u-x)du=\int\psi(x+y)\frac{y}{T-t}f_{T-t}(y)dy$
$=\mathbb{E}\left[\psi(x+Z)\frac{Z}{T-t}\right]=\mathbb{E}\left[\left(\psi(x+Z)-\psi(x)\right)\frac{Z}{T-t}\right]$
$\leq\frac{K}{T-t}\mathbb{E}\left[|Z|^{\alpha+1}\right]=KC_{\alpha}(T-t)^{\frac{\alpha+1}{2}-1},$
[…]
$\int_{0}^{T}|H_{s}|^{2}ds=\int_{0}^{T}|\partial_{x}u(t,W_{t})|^{2}dt\leq K^{2}C_{\alpha}^{2}\int_{0}^{T}(T-t)^{\alpha-1}dt\leq K^{2}C_{\alpha}^{2}T^{\alpha}.$
For general case, we can iterate the Step 1 backwards in time to obtain $L^{\infty}$ bound of $\int_{0}^{T}|H_{s}|^{2}ds$. ∎
We are now in a position to state the main result of this section. For the reader’s convenience, we first restate the main theorem below in a simplified setting and then generalize the statement. To complete the proof of the following theorem, it remains to extend the existence result from Malliavin differentiable terminal conditions to the general bounded case. Uniqueness then follows directly from the stability result established in Theorem 4.6.
@@
Let Assumptions $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{unif}})$} hold. In addition, assume that
$\xi=\phi(W_{t_{1}},W_{t_{2}},\cdots,W_{t_{n}})$
for a bounded function $\phi:(\mathbb{R}^{d})^{n}\to\mathbb{R}$ which is $\alpha$-Hölder continuous, that is, there exists $K>0$ and $\alpha\in(0,1]$ such that
$|\phi(x)-\phi(x^{\prime})|\leq K\max_{i=1,2,\cdots,n}|x_{i}-x^{\prime}_{i}|^{\alpha},$
for any $x,x^{\prime}\in(\mathbb{R}^{d})^{n}$. Then, there exists a unique solution $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the BSDE (2.6).
@@
Let $\eta_{m}$ be a mollifier on $(\mathbb{R}^{d})^{n}$ with compact support $([-\frac{1}{m},\frac{1}{m}]^{d})^{n}$. Consider $\phi^{m}(x):=\eta_{m}*\phi$, which is Lipschitz. Then $\xi^{m}:=\phi^{m}(W_{t_{1}},W_{t_{2}},\cdots,W_{t_{n}})$ has bounded Malliavin derivatives, and Lemma 5.1 yields a corresponding solution $(Y^{m},Z^{m})$. Let $H^{m}$ be the martingale representation integrand of $\xi^{m}$, that is,
$\xi^{m}=\mathbb{E}\xi^{m}+\int_{0}^{T}H^{m}_{s}dW_{s},$
and show that $\sup_{m}\int_{0}^{T}|H^{m}_{s}|^{2}ds\leq C$ almost surely for some $C$. Since $\phi^{m}$ has the same Hölder constant as $\phi$ for any $m$, we can apply Lemma 5.2 to obtain such a constant $C$. Finally, Theorem 4.6 provides the convergence of $(Y^{m},Z^{m})$ to $(Y,Z)$ in $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$. Uniqueness follows from Theorem 4.6 as well. ∎
Theorem 5.4.
Let Assumptions $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{sep}})$} hold. In addition, assume that there exists a function $\phi:C([0,T];\mathbb{R}^{d})\to\mathbb{R}$ such that $\xi=\phi(W_{[0,T]})$ and uniformly continuous in sup norm. Then, there exists a unique solution $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the BSDE (2.6).
Proof.
Let us consider the Moreau-Yosida approximation of $\phi$:
$\phi^{n}(\omega):=\inf_{y\in C([0,T];\mathbb{R}^{d})}\left\{\phi(y)+n\sup_{t\in[0,T]}\left|y(t)-\omega(t)\right|\right\}.$
Since $\inf_{y\in C([0,T];\mathbb{R}^{d})}\phi(y)\leq\phi^{n}(\omega)\leq\phi(\omega),$ $\phi^{n}$ is bounded. In addition, note that, for any $\omega,\omega^{\prime}$
$\phi^{n}(\omega)$ $=\inf\left\{\phi(y)+n\sup_{t\in[0,T]}\left|y(t)-\omega(t)\right|\right\}\leq\inf\left\{\phi(y)+n\sup_{t\in[0,T]}\left|y(t)-\omega^{\prime}(t)\right|\right\}+n\sup_{t\in[0,T]}\left|\omega^{\prime}(t)-\omega(t)\right|$
$\leq\phi^{n}(\omega^{\prime})+n\sup_{t\in[0,T]}\left|\omega^{\prime}(t)-\omega(t)\right|.$
Therefore, $\phi^{n}$ is Lipschitz in sup norm. Lastly, since $\phi$ is uniformly continuous, for $\varepsilon>0$, there exists $N$ such that $\sup_{t\in[0,T]}\left|y(t)-\omega(t)\right|<\frac{\varepsilon}{2n}$ implies $|\phi(y)-\phi(\omega)|\leq\varepsilon/2$ for all $n\geq N$. Then,
$|\phi^{n}(\omega)-\phi(\omega)|$ $=\left|\inf_{y\in C([0,T];\mathbb{R}^{d})}\left\{\phi(y)-\phi(\omega)+n\sup_{t\in[0,T]}\left|y(t)-\omega(t)\right|\right\}\right|<\varepsilon.$
Therefore, $\phi^{n}(\omega)$ converges to $\phi(\omega)$ uniformly in $\omega$.By Proposition 3.2 of Cheridito and Nam (2014), $\xi^{n}:=\phi^{n}(W)$ has bounded Malliavin derivatives. By Lemma 5.1, each approximating BSDE with terminal condition $\xi^{n}$ admits a solution $(Y^{n},Z^{n})\in\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$, satisfying